The annual profit for the firm depends upon the number of units produced. Specifically the function which describes the relationship between profit P (stated in dollars) and the number of units produced x is

P=-0.01x^2+5000x-25000

a) Determine the number of units x which will result in maximum profit.

b) What is the expected maximum profit?


1
Expert's answer
2021-01-25T00:06:13-0500

Given that

"P = -0.01x^{2} + 5000x - 25000"

Then we have that "\frac{dP}{dx} = -0.02x + 5000"

So when "\frac{dP}{dx} = 0"

We have that "-0.02x + 5000 = 0"

"0.02x = 5000"

"x = \frac{5000}{0.02}"

"x = 250000"

So the answer to question (a) is that 250000 will result in the maximum profit

b) So we substitute the value of "x = 250000" into "P = -0.01x^{2} + 5000x - 25000"

So we have that "P(250000) =" "-0.01(250000)^{2} + 5000(250000) - 25000"

"P(250000)= 624975000"

The expected maximum profit is 624975000